Circular No. 6314

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1994Y IN NGC 5371 P. Garnavich, Harvard-Smithsonian Center for Astrophysics; A. Nori-

ega-Crespo, Infrared Processing and Analysis Center; and A. Moro-Martin, Maria Mitchell Observatory, write: “Infrared photometry with the Apache Point 3.5-m telescope (+ GRIM-II array) shows that the peculiar type-II supernova 1994Y (IAUC 6058) has developed a strong infrared excess. Magnitude estimates for Jan. 14 are K = 15.6, H = 17.1, and J = 18.9. The excess is similar to that seen in SN 1980K (Dwek et al. 1983, Ap.J. 274, 168) and suggests dust formation or a strong infrared echo from an existing dust shell. The difference in distance between the host galaxies implies that, during this dust emission phase, SN 1994Y is intrinsically brighter than was SN 1980K by at least 3 magnitudes in the K band.”

GRO J1744–28 A. Żytkow, Institute of Astronomy, Cambridge; and M. Irwin, Royal

Greenwich Observatory, Cambridge, communicate: “The APM facility in Cambridge was used to analyze the area covering the position of GRO J1744–28 on glass copies of a U.K. Schmidt Telescope (UKST) photographic sky survey IIIa-J plate and a European Southern Observatory (ESO) R plate. For each plate, an ∼ 4′×4′ region centered on the VLA position (IAUC 6307) was measured. Our examination of the data clearly suggests a candidate for an optical counterpart that is within 1′′ of the VLA position. The estimated magnitude of this candidate on the ESO R plate is ∼ 19.5, whereas on the IIIa-J sky survey plate, it cannot be brighter than B_{j} ∼ 22, as it is close to the plate limit. The derived positions (equinox 2000.0) are: α = 17^{h}44^{m}36^{s}.92, δ = −28^{o}45′38′′.0 (UKST B_{j}); 36^{s}.78, 37′′.6 (position end figures; ESO R). The coordinates derived from the ESO R plate are more reliable, since the object is well above the plate limit (estimated errors ± 0′′.5); for the IIIa-J plate, the errors are ∼ 1′′. Thus, the candidate optical counterpart has an observed probable B_{j} − R > 2.5, equivalent to B − V > 1.9. In addition, visual inspection of the region on a UKST sky survey film copy in the I band (IV-N emulsion), shows the candidate optical counterpart of the VLA source to be significantly redder than neighboring objects, with estimated mag I = 17.5. If the above optical candidate proves to be related to GRO J1744–28, then it may be a new candidate for a class of neutron star/M star binaries similar to HD 154791.”

1996 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green